1. What is the H+ concentration of the soil when the pH is 4.5? What will the pH be if the H+ concentration increases 100-fold?

2)If the AEC of 100 g of a soil is 5 cmol (-) kg-1, what is its maximum capacity for the retention of Cl-?

3)If the pK of NH3 is 9.2, how much NH3 relative to NH4+ will be found at pH 6.5?

4)What is the maximum amount of Mg2+ that 50 g of a soil can hold if the soil has a CEC of 7 meq 100 g-1?

5)What is the surface area of 25 ug of spherical clay particles having a uniform diameter of 0.5 um and a particle density of 2.5 g cm-3? Hint: What are the surface area and volume of one sphere? You have to make some assumptions to solve this problem.
Answer #1

1. pH = -log(H+)

4.5=-log(H+)

-4.5=log(H+)

Then take antilog of -4.5 that is 10^-4.5

So H+ conc=3.1*10^-5

If we increase the conc. Of H+ 1 then conc . Will be 3.1*10^-3

Then ph= --log(H+)

pH = -log(0.0031)=- (-2.50)=2.50

If this solution really helps then give me a thumps up thank you